[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Aug 23, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Aug 23 09:48:01 UTC 2016


HSU Shift Log for Tue Aug 23 2016 05:47:54 GMT-0400 (EDT)

Shift: A

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None. Closed out all ATCF windows and re-opened them with no  
issues afterwards.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160823/28a4e806/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list